OpenStudy (anonymous):

Ben has four more than twice as many CDs as Jake. If they have a total of 31 CDs, how many CDs does Jake have? (1) 9 (3) 14 (2) 13 (4) 22

OpenStudy (anonymous):

B = the number of CDs owned by Ben J = the number of CSs owned by Jake Now you can write 2 different equations using J and B and the info given in the problem. Can you construct one of the equations?

OpenStudy (anonymous):

I know that i dont know how to get to the answer

OpenStudy (anonymous):

you can't just go straight for the answer, and nobody is going to just give you the answer here. work with me. I got you started. the total number of cds is 31. what equation, with the variables J and B, can you make using that total of 31?

OpenStudy (anonymous):

2J=31 and 4B=31?

OpenStudy (anonymous):

where did you get 2 and 4?

OpenStudy (anonymous):

okay continue

OpenStudy (anonymous):

B = the number of CDs owned by Ben J = the number of CSs owned by Jake The total number of their CDs combined is 31. Write an equation reflecting the above facts. You can do it.

OpenStudy (anonymous):

2B+4=31

OpenStudy (anonymous):

When you see "total" think "add everything up" Add Ben's CDs and Jake's CDs together and the result is 31 CDs

OpenStudy (anonymous):

In English: Ben's CDs plus Jake's CDs equal 31 CDs In Algebra: B + J = 31

OpenStudy (anonymous):

2J+4B=31

OpenStudy (anonymous):

I have no idea what 2J + 4B = 31. I just gave you the equation I was looking for: B + J = 31

OpenStudy (anonymous):

NOW we can construct an equation reflecting that first sentence: "Ben has four more than twice as many CDs as Jake" Ben CDs = 4 + (2 x Jake's CDs)

OpenStudy (anonymous):

B = 4+ 2J

OpenStudy (anonymous):

2 equations, 2 unknowns: B + J = 31 B = 4 + 2J Use substitution to solve for B and J Good luck.

OpenStudy (anonymous):

so j equals 9

OpenStudy (anonymous):

thank you